# AutoBencher: Creating Salient, Novel, Difficult Datasets for Language Models

### v1

@@ 4.1 Generating Datasets with Privileged Information
Figure 2: How the model $\texttt{LM}_{\text{gen}}$ uses privileged information to create (question, answer) examples.
To improve the difficulty of the generated questions and the correctness of the generated answers, we augment $\texttt{LM}_{\text{gen}}$ with privileged information that we do not provide to the LMs that we are evaluating (which we will refer to as the test-taker LMs to avoid confusion). This privileged information ensures answer correctness by grounding the generated answers in a reliable source and creates an information imbalance between the example creation model $\texttt{LM}_{\text{gen}}$ and test-taker LMs, which simplifies the task for $\texttt{LM}_{\text{gen}}$ and enables it to create questions that are more difficult than possible with its base capabilities. We next detail the privileged information we provide in three domains: knowledge intensive, multilingual, and mathematics. Figure 2 illustrates how this information is used in each domain.
Knowledge-intensive.
We augment $\texttt{LM}_{\text{gen}}$ with a set of relevant documents (e.g., Wikipedia). Specifically, to create a knowledge-intensive question on topic $c$, we first retrieve the set of most relevant articles by querying $c$ in the Wikipedia Search API. Then, we prompt $\texttt{LM}_{\text{gen}}$ to jointly generate (question, answer) pairs conditioned on the retrieved articles such that the answer can be verified by the document and the question is answerable without the document. We enforce this by checking the lexical overlap between the answers and the retrieved articles.
Multilingual.
We augment $\texttt{LM}_{\text{gen}}$ with a translation system. Since models tend to have better reasoning capabilities in English than in other languages, we generate (question, answer) pairs in other languages by first generating the example in English via the knowledge-intensive question procedure above. Then, we translate the question and answer to the target language.
Mathematics.
We augment $\texttt{LM}_{\text{gen}}$ with Python math libraries (e.g., sympy, scipy, numpy). To ensure that the answers are correct, we prompt $\texttt{LM}_{\text{gen}}$ to generate questions along with Python code to compute their answers and use the execution result as the answer.
@@ 4.2 Proposing Topics with Adaptive Search
When we use $\texttt{LM}_{\text{gen}}$ to propose evaluation topics, a key challenge is that $\texttt{LM}_{\text{gen}}$ does not have information about what topics might be difficult for LMs. To address this, we propose adaptive search, an iterative approach that collects accuracy information in each iteration to inform topic proposals in subsequent iterations. We keep track of a trajectory $\mathcal{H}$, represented as a sequence of (topic, accuracy) pairs. As we run more iterations, $\mathcal{H}$ accumulates more (topic, accuracy) pairs, and forms a better belief about what topics are likely be difficult (see Figure 1 right).
Specifically, we follow lines 1 to 7 in Algorithm 1. In each iteration, AutoBencher proposes $K$ topics conditioned on the trajectory $\mathcal{H}$ collected from previous iterations (line 3), where we specifically prompt to ask for difficult topics (i.e., low accuracy). We filter out non-salient topics (line 4) and instantiate each remaining topic as a dataset via the procedure described in § 4.1 (line 5). Then, we compute the accuracy of a test-taker model $\texttt{LM}_{\text{test}}$ on each dataset as a measure of difficulty (line 6). Finally, we feed all proposed (topic, accuracy) pairs to subsequent iterations to increase difficulty (lines 7). We choose to only evaluate the accuracy of a single strong test-taker model, rather than evaluating all models $\mathcal{M}$ for computational efficiency.
Our adaptive search procedure does not take Novelty or Sep into account, since these two quantities also require expensively evaluating all models $\mathcal{M}$. Instead, we optimize these desiderata in a final re-ranking step, which considers all of the metrics together.
Require: $\mathcal{M}$, $\texttt{LM}_{\text{gen}}$, $\texttt{LM}_{\text{test}}$, domain $d$, max iterations $N$, hyperparameter $\beta_{1}$, $\beta_{2}$ 1:Initialize previously-proposed topics $\mathcal{H}=\varnothing$ 2:for maximimum number of iteration $N$ times do 3: Propose topics conditioned on prev. topics $c_{1},\ldots,c_{K}\sim\texttt{LM}_{\text{gen}}(\cdot\mid\mathcal{H})$ 4: Filter out to keep only the salient topics with $c\in\mathcal{S}$ 5: Generate small dataset $\mathcal{D}_{c_{k}}$ for each remaining topic by prompting $\texttt{LM}_{\text{gen}}$ 6: Compute the test-taker model accuracy on each dataset $\texttt{acc}(\texttt{LM}_{\text{test}},\mathcal{D}_{c_{k}})$ 7: Update previously proposed topics =$\mathcal{H}=\mathcal{H}\cup\{(c_{k},\texttt{acc}(\texttt{LM}_{\text{test}},\mathcal{D}_{c_{k}}))\}$ 8:Extract set of all proposed topics $\mathcal{P}=\{c:(c,\texttt{acc}(\texttt{LM}_{\text{test}},\mathcal{D}_{c}))\in\mathcal{H}\}$ 9:Compute the search objective $\mathcal{J}(c)$ on all proposed topics $c\in\mathcal{P}$ 10:Select the topic with the highest objective value $c^{*}=\arg\max_{c\in\mathcal{P}}\mathcal{J}(c)$ 11:Generate large dataset $\mathcal{D}_{c^{*}}$ by prompting $\texttt{LM}_{\text{gen}}$ on topic $c^{*}$ 12:return Chosen topic $c^{*}$ and corresponding dataset $\mathcal{D}_{c^{*}}$
@@
4.3 Metric Attribution
Overall, the entire algorithm behind AutoBencher is summarized in Algorithm 1. In lines 1 to 7, we perform the adaptive search described above. Afterward, we re-rank the proposed topics via the full search objective $\mathcal{J}(c)$: We compute the objective for each proposed topic (line 9) and construct a dataset on the topic with the highest objective value (lines 10–12).
This algorithm optimizes for all the desiderata of novelty, salience, and difficulty. salience is met via the filtering step in line 4, which filters out all non-salient topics. Difficulty is approximately optimized via the adaptive search procedure (lines 1–7) that steers toward producing more difficult topics. Sep and Novelty are optimized in the final re-ranking step (lines 9–10), where we compute and select based on the full objective function $\mathcal{J}$.
@@ 5.1 Baselines and Metrics
We evaluate AutoBencher on five domains: mathematics, multilingual, and knowledge-intensive QA on history, economy, and science.
Baselines.
We compare benchmarks generated by AutoBencher with human constructed benchmarks, which we group collectively under HumanBench. For knowledge-intensive domains, we compare with the subject datasets in MMLU [11]. These datasets consist of multiple-choice questions on academic subjects ranging from high school to professional level. We compare with 4 history subjects (e.g., high school world history), 4 economy subjects (e.g., econometrics), and 7 science subjects (e.g., college physics). See the complete list in Appendix C.
For mathematics, we compare with $7$ datasets from the Mathematics Dataset [31] [^3] https://github.com/google-deepmind/mathematics_dataset, which covers basic math capabilities: algebra, arithmetic, calculus, probability, comparison, measurement, numbers. For multilingualism, we compare with XOR QA [2], a multilingual question-answering dataset covering 7 diverse languages. We compare with the test set, split by language into 7 datasets.
Models.
[…]
Metrics.
We evaluate on the four metrics: Novelty (Novel), Sep (Sep), Difficulty (Diff), and salience as defined in § 3. For calculating Novelty, we set $\mathbf{D}_{\text{prev}}$ as the aggregate of all datasets in HumanBench.
@@ 5.2 AutoBencher Hyperparameters and Costs
Hyperparameters. AutoBencher uses gpt-4-0125-preview [24] as $\texttt{LM}_{\text{gen}}$ to propose topics and generate the datasets. For the test-taker model $\texttt{LM}_{\text{test}}$, we select gpt-3.5-turbo-0613 [23], Mixtral-8x7B and Mistral-7B, to cover different levels of model accuracies. In AutoBencher, we perform $N=8$ iterations of adaptive search, each proposing $k=5$ topics. We generate $|\mathcal{D}_{c}|=50$ examples per topic. In total, we explore 2000 questions across 40 topics in one round of AutoBencher.
Costs. Each run of the AutoBencher agent uses around 750K tokens, which costs around \$15. Among them, 43K tokens are used for proposing topics, 576K tokens are used for constructing datasets, and 147K for evaluation.
@@ 6 Main Results
We find that AutoBencher successfully constructs datasets that satisfy the three desiderata. We first report the novelty and difficulty scores in § 6.1. Then we provide the list of discovered topics and qualitative examples of questions generated by AutoBencher in § 6.2. Finally, we conduct a human study that verifies the salience, novelty, and difficulty of the discovered topics in § 6.3.
6.1 Novelty and Difficulty
Recall that our definition of novelty measures the rank correlation between models’ accuracies on one dataset with their accuracies on all other datasets. A lower correlation indicates more novel performance trends. We find that datasets constructed by AutoBencher are significantly more novel than existing human-constructed datasets, reducing the rank correlation by 27%. Moreover, AutoBencher datasets also exhibit 22% greater difficulty (Diff) and higher separation (Sep) between models, increasing the accuracy gaps between existing models by 1%, on average. These improvements hold across all domains, as shown in Table 1.
We evaluate the impact of adaptive search on novelty and difficulty by ablating it in AutoBench-AS. Rather than conditioning on the (topic, accuracy) pairs of previously proposed topics, we simply prompt $\texttt{LM}_{\text{gen}}$ to propose salient, difficult, and diverse topics. Table 1 (top) shows that AutoBench-AS obtains lower novelty and difficulty scores than full AutoBencher, but still outperforms the human-constructed datasets in all metrics. This is likely because adaptive search only affects the quality of the proposal distribution, and AutoBench-AS still accounts for novelty and difficulty via final re-ranking on the objective function.
Table 1: Comparison between AutoBencher and prior human-constructed datasets (HumanBench) on novelty, separation (Sep), and difficulty (Diff). Higher numbers are better for all metrics. AutoBencher constructs datasets that are significantly more novel and difficult over human-constructed datasets. Ablating the adaptive search component impacts all metrics, particularly difficulty.
@@ 6.2 Qualitative Examples
To qualitatively analyze the results of AutoBencher, we list the topics that are selected by AutoBencher along with the number of Wikipedia page views of that topic and $3$ random (question, answer) pairs from each constructed dataset in Table 3. We find that the generated questions and answers are generally of high quality: they are primarily factually correct and self-contained, though some questions are too open-ended or have dubious answers, e.g., “What is the significance of blockchain technology in environmental science?” Below, we also discuss the novel performance trends AutoBencher reveals in four topics. See Appendix D for the full ranking details.
Fordism is an important economic concept on mass production. Gemini Pro drops from an average rank of $6$ on existing economics datasets to $16$. In contrast, OpenChat-3.5 performs quite well. It jumps from an average economics rank of 8 to 2 on this dataset, nearly matching Claude 3 Sonnet.
Secret Society is a historical topic, and exhibits better-than-expected GPT-3.5 Turbo accuracy, rising from rank $7$ to $3$. Gemini Pro performs worse than expected, dropping from rank $6$ to $16$.
Economic impact of COVID-19 reveals much better-than-expected accuracy by Mistral-7B and OpenAGI-7B, which nearly match GPT-4 Turbo, perhaps due to the wealth of COVID-19 data.

### v2

@@ 4.1 Generating Datasets with Privileged Information
Figure 2: How the model $\texttt{LM}_{\text{evaluator}}$ uses privileged information to create (question, answer) examples.
To improve the difficulty of the generated questions and the correctness of the generated answers, we augment $\texttt{LM}_{\text{evaluator}}$ with privileged information (denoted as $\mathcal{I}$). The privileged information (e.g., Wikipedia articles in Figure 2) is only available to the evaluation LM, improving correctness by grounding the generated answers in a reliable source. It’s not provided to the candidate LMs, which creates an information asymmetry between the evaluator LM and candidate LMs. Specifically, the evaluator LM generates (question, answer) pairs: $(q,a)\sim\texttt{LM}_{\text{evaluator}}(\mathcal{I},c)$, and the candidate LMs answer these questions: $\hat{a}\sim\texttt{LM}_{\text{candidate}}(q)$. Augmented with privileged information simplifies the task for $\texttt{LM}_{\text{evaluator}}$ and enables it to create questions that are more difficult than possible with its base capabilities. Figure 2 illustrates how this information is used in each domain.
We next detail the privileged information we provide in three domains: knowledge intensive, multilingual, and mathematics. In Appendix E, we discuss more examples of privileges based on the compute and problem structure.
Knowledge-intensive domains. We augment $\texttt{LM}_{\text{evaluator}}$ with a set of relevant documents (i.e., $\mathcal{I}$ is relevant Wikipedia articles). Specifically, to create knowledge-intensive questions relevant to the natural language description $c$, we first retrieve the set of most relevant articles by querying $c$ in the Wikipedia Search API. Then, we prompt $\texttt{LM}_{\text{evaluator}}$ to jointly generate (question, answer) pairs conditioned on the retrieved articles. Concretely, we want the question to be answerable without the document (i.e., answerable by the candidate LMs without the privileged information) and the generated answer to be verified by the document (i.e. correctness).
Multilingual domains. We augment $\texttt{LM}_{\text{evaluator}}$ with a translation system (i.e., $\mathcal{I}$ is a multilingual LM prompted to translate text from English to a target language). Since models tend to have better reasoning capabilities in English than in other languages, we generate (question, answer) pairs by first generating the example in English via the knowledge-intensive question procedure above. Then, we translate the question and answer to the target language.
Math domains. We augment $\texttt{LM}_{\text{evaluator}}$ with Python math libraries (e.g., $\mathcal{I}$ is Python libraries like sympy, scipy, numpy). To ensure that the answers are correct, we prompt $\texttt{LM}_{\text{evaluator}}$ to generate questions along with Python code to compute their answers and use the execution result as the answer. The candidate LMs need to answer the math questions directly, without calling Python libraries.
Safety domains. We do not use privileged information in the safety domain. Privileged information is not needed to generate correct answers to harmful requests, because the correct responses are always to abstain (e.g., “I can’t assist with that. ”). Therefore, we set $\mathcal{I}=\emptyset$ and prompt the $\texttt{LM}_{\text{evaluator}}$ to generate the harmful requests $q\sim\texttt{LM}_{\text{evaluator}}(\emptyset,c)$.
@@ 4.2 Proposing Topics with Adaptive Search
[…]
We present the full AutoBencher algorithm in Algorithm 1. Adaptive search refers to lines 1 to 7 in Algorithm 1. In each iteration, AutoBencher proposes $K$ descriptions conditioned on the trajectory $\mathcal{H}$ collected from previous iterations (line 3), where we specifically prompt to ask for dataset descriptions that elicit low model accuracies. We filter out non-salient descriptions (line 4) and construct a dataset from each remaining description, augmented with privileged information (line 5; § 4.1). Then, we compute the accuracy of a candidate LM on each dataset as a measure of difficulty (line 6). Finally, we feed all proposed (description, accuracy) pairs to the next iteration (lines 7).
Our adaptive search procedure does not take novelty or separability into account, since these two quantities require evaluating all models $\mathcal{M}$. Instead, we take these factors into account in a final re-ranking step via the full search objective $\mathcal{J}(c)$: We compute the objective for each proposed dataset description (line 9) and output a dataset on the description that achieves the highest objective value (lines 10–12).
Require: a evaluator language model $\texttt{LM}_{\text{evaluator}}$, a candidate language model $\texttt{LM}_{\text{candidate}}$, domain $d$, max iterations $N$, number of dataset descriptions per iteration $K$
1:Initialize previously-proposed dataset descriptions $\mathcal{H}=\varnothing$
2:for maximimum number of iteration $N$ times do
3: Propose dataset descriptions conditioned on prev. descriptions $c_{1},\ldots,c_{K}\sim\texttt{LM}_{\text{evaluator}}(\cdot\mid\mathcal{H})$
4: Filter out to keep only the salient (or harmful) descriptions with $c\in\mathcal{S}$
5:for $c$ in the remaining descriptions do
6:  Generate small dataset $\mathcal{D}_{c}$ for each by prompting $\texttt{LM}_{\text{evaluator}}$ with privileged information.
7:  Compute the test-taker model accuracy on each dataset $\texttt{acc}(\texttt{LM}_{\text{candidate}},\mathcal{D}_{c})$
8:  Update previously proposed topics =$\mathcal{H}=\mathcal{H}\cup\{(c,\texttt{acc}(\texttt{LM}_{\text{candidate}},\mathcal{D}_{c}))\}$
9:Extract set of all proposed descriptions $\mathcal{P}=\{c:(c,\texttt{acc}(\texttt{LM}_{\text{candidate}},\mathcal{D}_{c}))\in\mathcal{H}\}$
10:Compute the search objective $\mathcal{J}(c)$ on all proposed description $c\in\mathcal{P}$
11:Select the description with the highest objective value $c^{*}=\arg\max_{c\in\mathcal{P}}\mathcal{J}(c)$
12:Generate large dataset $\mathcal{D}_{c^{*}}$ by prompting $\texttt{LM}_{\text{evaluator}}$ on description $c^{*}$
13:return chosen dataset description $c^{*}$ and corresponding dataset $\mathcal{D}_{c^{*}}$
@@ 5 Experimental Setup
We evaluate AutoBencher for the capabilities and safety. Within the capabilities settings, we consider six domains: mathematics, multilingualism, history, economy, and science.
@@ 5.1 Baselines and Metrics
Baselines. For the capability settings, We compare benchmarks generated by AutoBencher with human-constructed benchmarks (denoted as HumanBench). For knowledge-intensive domains, HumanBench contains datasets in MMLU (Hendrycks et al., 2021), including 4 history subjects (e.g., high school world history), 4 economy subjects (e.g., econometrics), and 7 science subjects (e.g., college physics). See the complete list in Appendix C. For mathematics, HumanBench contains $7$ datasets from the Mathematics Dataset (Saxton et al., 2019) [^2] https://github.com/google-deepmind/mathematics_dataset, which covers basic math capabilities: algebra, arithmetic, calculus, probability, comparison, measurement, numbers. For multilinguality, we compare with XOR QA (Asai et al., 2021), a multilingual question-answering dataset covering 7 diverse languages. We compare with the test set, split by language into 7 datasets.
For the safety setting, we compare with XSTest (Röttger et al., 2024) and HarmBench (Mazeika et al., 2024), which are popular safety datasets that evaluate whether a model can accurately reject harmful requests.
Models. We evaluate on the model family of GPT-4, GPT-3.5, Claude-3, Claude-2, Mixtral, Mistral, Gemini, LLaMA-2, LLaMA-3 and LLaMA’s finetuning derivatives. See Appendix D for the full model list.
Metrics. For the capability setting, we evaluate on the three metrics: Novelty (Novel), separability (Sep), and Difficulty (Diff) as defined in § 3. For calculating Novelty, we set $\mathbf{D}_{\text{prev}}$ as the aggregate of all datasets in HumanBench.
For the safety setting, we report the average attack success rate (ASR) of the datasets, as defined in § 3.
@@ 5.2 AutoBencher Hyperparameters and Costs
Hyperparameters. AutoBencher uses gpt-4-0125-preview (OpenAI, 2023) as $\texttt{LM}_{\text{evaluator}}$ (at temperature 0) to propose topics and generate the datasets. To construct a capability dataset, we perform $N=8$ iterations of adaptive search, each proposing $K=5$ descriptions, and we generate $|\mathcal{D}_{c}|=50$ examples per description. In the optimization objective, $\beta_{1}=1$ and $\beta_{2}=10$ are chosen so that the three terms have similar scales. To construct a safety dataset, we perform $N=10$ iteration of adaptive search, each proposing $K=10$ descriptions, and we generate $10$ examples for each description. For knowledge-intensive and multilingual questions, a dataset description is considered salient if the corresponding Wikipedia article has 500K+ views. For math and safety domains, we manually judge the salience of the dataset descriptions and remove the non-salient or non-harmful ones. See more details in Appendix D.
Costs. Each run of the AutoBencher agent uses around 750K tokens, which costs around \$15. Among them, 43K tokens are used for proposing topics, 576K tokens are used for constructing datasets, and 147K for evaluating the candidate LMs.
@@ 6 Main Results
We find that AutoBencher successfully constructs datasets that achieves our declared desiderata. We first report the novelty, difficulty, and separability scores for the capability datasets in § 6.1. Then we report the attack success rate of our safety datasets in § 6.2. We provide the list of discovered dataset descriptions and qualitative examples of questions generated by AutoBencher in § 6.3. Finally, we conduct human evaluation to verify the correctness and salience of AutoBencher datasets in § 6.4.
6.1 Capability Settings: Novelty, Difficulty, Separability
Recall that we define novelty to measure the rank correlation between models’ accuracies on one dataset with their accuracies on all other datasets [^3] See Appendix D for a full list of models we evaluate. . A lower correlation indicates more novel performance trends. We find that datasets constructed by AutoBencher are significantly more novel than existing human-constructed datasets, reducing the rank correlation by 27%. Moreover, AutoBencher datasets also exhibit 22% greater difficulty (Diff) and higher separation (Sep) between models, increasing the accuracy gaps between existing models by 1%, on average. These improvements hold across all domains, as shown in Table 1.
We evaluate the impact of adaptive search on novelty and difficulty by ablating it in AutoBench-AS. Rather than conditioning on the (description, accuracy) pairs of previously proposed topics, we simply prompt $\texttt{LM}_{\text{evaluator}}$ to propose salient, difficult, and diverse topics. Table 1 (top) shows that AutoBench-AS obtains lower novelty and difficulty scores than full AutoBencher, but still outperforms the human-constructed datasets in all metrics. This is likely because adaptive search only affects the quality of the proposal distribution, and AutoBench-AS still accounts for novelty and difficulty via final re-ranking on the objective function.
Table 1: Comparison between AutoBencher and prior human-constructed datasets (HumanBench) on novelty (Novel), separation (Sep), and difficulty (Diff). Higher numbers are better for all metrics. AutoBencher constructs datasets that are significantly more novel and difficult over human-constructed datasets. Ablating the adaptive search component (AutoBench-AS) degrades all metrics, particularly difficulty.
@@ 6.1 Capability Settings: Novelty, Difficulty, Separability
ASR
XSTest 0.08
HarmBench 0.28
AutoBench $\mathbf{0.38}$
HarmBench GCG-T 0.45
6.2 The Safety Setting: Attack Success Rate
We find that the AutoBencher dataset reveals more safety vulnerabilities than existing human-constructed datasets. As shown in Table 1, AutoBencher improves the attack success rate (ASR) by 20% on average. This suggests that our approach successfully discovers unsafe questions that existing models fail to defend against. AutoBencher does not outperform direct adversarial attacks like GCG [^4] The GCG prompts would not satisfy the harmfulness desiderata because it contains random tokens that are not fluent., because AutoBencher does not optimize for each request; instead, it searches for systematic categories of failures. One can imagine applying GCG on AutoBencher-generated requests to further enhance the ASR.
@@ 6.3 Qualitative Examples
To qualitatively analyze the results of AutoBencher, we provide some dataset descriptions selected by AutoBencher and randomly sampled (question, answer) pairs from each dataset in Table 2 and Table 13. For the capability datasets, the generated questions are salient and the generated answers are factually correct. For the safety datasets, the generated questions are unsafe, and the candidate LMs fail to abstain. Below, we discuss 3 novel performance trends and 2 safety risks revealed by AutoBencher and report models’ leaderboard rankings.
Fordism is an important economic concept on mass production. Gemini Pro drops from an average rank of $6$ on economic HumanBench  to $16$ on Fordism. In contrast, OpenChat-3.5 performs quite well. It jumps from an average rank of 8 to 2 on this dataset, nearly matching Claude 3 Sonnet.
Secret Society is a historical topic, and exhibits better-than-expected GPT-3.5 Turbo accuracy, rising from the average rank of $7$ to $3$. Gemini Pro performs worse than expected, dropping from rank $6$ to $16$.